# Privacy Technologies for Financial Intelligence

### v1

@@ 5 Discussion and Conclusion
From the practical perspective of being able to do something about complex financial crimes today, we advocate for the integration of privacy-preserving technologies within financial intelligence tools, not just as a mitigation strategy while waiting for legal experts to navigate the complex AML/CTF reform landscape [Bug18], but to inform those AML/CTF reform discussions. This approach offers several benefits. First, privacy-preserving technologies can be swiftly implemented, allowing financial institutions and financial intelligence units to address data privacy concerns while fulfilling AML/CTF obligations. Second, adopting privacy-preserving methods helps institutions build trust with customers and stakeholders by demonstrating a commitment to both data protection and effective financial crime prevention. Third, privacy-preserving solutions offer a flexible framework that can adapt to evolving regulatory requirements, helping institutions stay compliant as laws and regulations are updated.
@@
We hope this paper enhances public understanding of privacy-preserving financial intelligence through the various AML/CTF use cases we explore. The use cases highlight the evolving nature of financial crimes and the necessity for ongoing innovation in privacy technologies to address emerging threats effectively. Having said that, for financial institutions and regulatory bodies, adopting privacy technologies involves careful consideration of both technological and operational aspects. Effective implementation strategies must balance the trade offs between data privacy and the quality and efficiency of analysis. While these technologies address immediate challenges, they are not a substitute for comprehensive regulatory reforms. Continuous refinement of regulations is essential to align financial crime prevention efforts with data privacy goals.
@@ Acknowledgements
We would like to thank the many former colleagues in AUSTRAC and financial crime units in major Australian financial institutions who spent many hours teaching us the intricacies of the financial intelligence business. We are also grateful to many of them for continuing to play an important role to keep our financial system and our country safe.

### v2

@@ 6 Discussion
From the practical perspective of being able to do something about complex financial crimes today, we advocate for the integration of privacy-preserving technologies within financial intelligence tools, not just as a mitigation strategy while waiting for legal experts to navigate the complex AML/CTF reform landscape [11], but also to inform those AML/CTF reform discussions. This approach offers several benefits. First, privacy-preserving technologies can be swiftly implemented, allowing financial institutions and financial intelligence units to address data privacy concerns while fulfilling AML/CTF obligations. Second, adopting privacy-preserving methods helps institutions build trust with customers and stakeholders by demonstrating a commitment to both data protection and effective financial crime prevention. Finally, privacy-preserving solutions offer a flexible framework that can adapt to evolving regulatory requirements, helping institutions stay compliant as laws and regulations are updated to address new financial technologies and emerging criminal threats.
Trade-offs and Practical Considerations
In practical financial intelligence applications, selecting the most appropriate privacy techniques is crucial for ensuring data security, regulatory compliance, and maintaining user trust at sustainable costs. There are often trade-offs involved in such choices. With the increasing industry focus on responsible AI practices [20, 33], there is also an emerging need to understand the possible interactions between privacy requirements and related issues like explainability and fairness [5, 49, 60]. Scholars and practitioners must look ahead to how AI itself may transform AML/CTF, potentially enabling new, automated, or hybrid methods especially when combined with emerging technologies such as decentralised autonomous organisations [30]. Strategic foresight and scenario-based methodologies may therefore be essential to anticipate and prepare for the future evolution of AI-enabled AML/CTF. To inform those broader issues, we offer here a few observations on how privacy-utility trade-offs are usually handled for the different privacy technologies.
First of all, it is worth noting that the use of homomorphic encryption (HE) and secure multiparty computation (MPC) in financial-intelligence algorithms like private set intersection, graph pattern matching, and statistical model learning do not usually result in any change to the output of those algorithms. This is because HE and MPC allow an algorithm to be evaluated in an encrypted space but they do not otherwise change the nature of a given algorithm. The primary consideration when using HE and MPC is thus whether the specific algorithm and parameters chosen provide, at reasonable organisational and monetary costs, sufficient provable security against adversaries with access to potentially large computational resources and arbitrary background and side information.
In contrast, the use of differential privacy (DP) and federated learning (FL) to privatise a financial-intelligence algorithm can result in actual changes to the output of the algorithm. There are rich and well-established frameworks [23, 35, 47] to study trade-offs between privacy and utility when using DP, and these frameworks work well in conjunction with standard techniques for model evaluation and selection [51] in both federated learning and more typical setups. Addressing these questions will be central to future research in this field.
7 Conclusion
@@
We hope this paper enhances public understanding of privacy-preserving financial intelligence through the various AML/CTF use cases we explore. The use cases highlight the evolving nature of financial crimes and the necessity for ongoing innovation in privacy technologies to address emerging threats effectively. Having said that, for financial institutions and regulatory bodies, adopting privacy technologies involves careful consideration of both technological and operational aspects. Effective implementation strategies must balance the trade-offs between data privacy and the quality and efficiency of analysis. While these technologies address immediate challenges, they are not a substitute for comprehensive regulatory reforms. Further and continuous refinement of regulations is essential to align financial crime prevention efforts with data privacy goals.
@@ Acknowledgements
We thank the many former colleagues in AUSTRAC and financial crime units in major Australian financial institutions who spent many hours teaching us the intricacies of the financial intelligence business. We are grateful to many of them for continuing to play an important role to keep our financial system and our country safe.